A reaction between tungsten (VI) oxide (WO3) and hydrogen gas produces tungsten and water. What is the percentage yield of tungsten if 56.9 g of tungsten oxide yields 41.4 g of tungsten?


1
Expert's answer
2021-03-10T18:04:30-0500

WO3 + 3H2---> W + 3H2O

Mass of WO3= 56.9g

Mass of W= 41.4g

Molar mass of WO3= 232g/mol

Molar mass of W= 184g/mol

From the balanced equation,

234g of WO3 gives 184g of W

56.9g of WO3 should give 184/232 x 56.9= 45.1g

Now, since only 41.4g of W is produced,

Percentage yield= 41.4/45.1 x 100 = 91.8%